OpenStudy (anonymous):

the answer i get in the back of my question book is the 2square root 13, the answer i get is square root 56, can someone tell me how it is that i'm suppose to obtain the description of the value, and i'm not aloud to use a calculator.

OpenStudy (phi):

i get is square root 56, what problem are you working on? if you got sqrt(52) you could say sqrt(13*4)= 2*sqrt(13)

OpenStudy (anonymous):

the question says a balloon is rising at 4m/s and the wind is blowing at rate of 6m/s parallel to the ground, find its resulting speed and its angle of point of descent

OpenStudy (phi):

yes, your book is correct.

OpenStudy (phi):

how did you sqrt(56)?

OpenStudy (anonymous):

4^2+6^2=56

OpenStudy (anonymous):

they both amount to the same value

OpenStudy (phi):

you added wrong. check again

OpenStudy (anonymous):

sorry thats 52 squared

OpenStudy (phi):

now factor 52 into 2*2*13 so sqrt(52) becomes 2*sqrt(13)

OpenStudy (anonymous):

whats a fast way to obtain those factors of this value

OpenStudy (phi):

learn some rules: if a number is even, the divisible by 2 sum of digits div by 3, then div by 3 ex: 126= 1+2+6=9/3 so 126/3 last 2 digits div by 4, then div by 4 end in 0 or 5, div by 5 div by 2 and 3, then div 6 sum of digits div by 9, then div by 9 plus, memorize the prime numbers, as many up to 100 as you can

OpenStudy (anonymous):

ok its been a while since ive had to do this but that jolts the memory a bit. noe i've got another question regarding vectors, do you want that one here or should i post it

OpenStudy (phi):

the angle